Question #222307

The number of calls coming into an hotels reservation centre is poisson random variable with mean 3

i) Find the probability that no calls come in a given one minute period

ii)Assume that the number of calls arriving in two different minutes are independent. Find the probability that at least two calls will arrive in a given two minute period.


Expert's answer

i) Let X=X= the number of calls arriving  in a given one minute period: X∼Po(λt)X\sim Po(\lambda t)

Given λ=3,t=1\lambda=3, t=1

λt=3(1)=3\lambda t=3(1)=3


P(X=0)=e−3⋅300!=e−3≈0.049787P(X=0)=\dfrac{e^{-3}\cdot3^0}{0!}=e^{-3}\approx0.049787


ii) Let X=X= the number of calls arriving in a given two minute period: X∼Po(λt)X\sim Po(\lambda t)

Given λ=3,t=2\lambda=3, t=2

λt=3(2)=6\lambda t=3(2)=6


P(X≥2)=1−P(X=0)−P(X=1)P(X\geq 2)=1-P(X=0)-P(X=1)




=1−e−6⋅600!−e−6⋅611!=1−7e−6≈0.982649=1-\dfrac{e^{-6}\cdot6^0}{0!}-\dfrac{e^{-6}\cdot6^1}{1!}=1-7e^{-6}\approx0.982649


LATEST TUTORIALS
APPROVED BY CLIENTS